# A Quarter of a Century of Neuromorphic Architectures on FPGAs - an Overview

### v2

@@ Abstract.
journal: CSURcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviews
1. Introduction
@@
Computers and their performance improvement have historically relied on two pillars: Moore’s law (Schaller, 1997) and Dennard’s scaling (Bohr, 2007). The former was formulated by Gordon Moore in the 1960s and states that the number of transistors per unit area double every 24 months, primarily due to improvements in lithography techniques. The latter was introduced by Richard Dennard, who found that as transistors grow smaller, the energy consumption reduces, resulting in constant power density. These two observations worked in tandem, allowing us to build more complex high-performing computers (Moore’s law) without them consuming needlessly large amounts of energy (Dennard’s scaling). Unfortunately, Dennard’s scaling ended in 2004-2005 (Dennard et al., 2018), and we are currently witnessing the slow-down of Moore’s law (Schaller, 1997) as well as the emergence of Dark Silicon (Esmaeilzadeh et al., 2011) without any in-place replacement technology at hand (Shalf, 2020). Instead, researchers and industry are looking for alternative ways of computing in hope of continuing the trend in computer performance we have seen so far. These technologies are collectively called post-Moore technologies (Leiserson et al., 2020; Shalf, 2020) and include quantum computers (Gyongyosi and Imre, 2019), adiabatic reversible logic (Bommi and Selvakumar, 2018), and analog computing (MacLennan, 2007). Two post-Moore technologies, however, are perhaps the most salient alternatives: Neuromorphic– (Schuman et al., 2017) and Reconfigurable–computing (Podobas et al., 2020; Kuon et al., 2008).
@@
Introduced in the 1980s by Carver Mead, neuromorphic systems (Mead, 1990) are computers built to replicate – to various extent – the impressive computing capabilities of the biological brain. Initially designed to replicate brain circuitry using analog electronic components, neuromorphic systems today are implemented in different technologies, ranging from digitally synchronous (e.g., Intel Loihi (Davies et al., 2018) or IBM TrueNorth (Akopyan et al., 2015)), mixed analog/digital (e.g., BrainDrop (Neckar et al., 2018)), fully analog (e.g. (Indiveri et al., 2011)), to even memristor- or photonics-based systems (Li et al., 2018; Shastri et al., 2021). Regardless of how they are implemented, they share two fundamental traits: (i) they communicate through discrete events called spikes and (ii) they are programmed in a non-imperative way, generally by encoding the problem to be solved through a spiking neural network (SNN) (Maass, 1997). Solving a problem using a neuromorphic system, often results in a significantly more energy-efficient solution in several application domains (Davies et al., 2021) as opposed to a traditional von Neumann system (Nikhil, 1989). A large number of functionalities of neuromorphic systems are driven by discoveries and innovation in neuroscience (e.g., Spike-Timing Dependent Plasticity (STDP) (Song et al., 2000), Bayesian Confidence Propagation Neural Network (BCPNN) (Ravichandran et al., 2024), or Eligibility Propagation (E-prop) (Bellec et al., 2020)), which has motivated many researchers to abandon non-reconfigurable and expensive application-specific integrated circuits (ASICs) and leverage the reconfigurable Field-Programmable Gate Arrays (FPGAs) to keep up with the fast pace of neuroscientific research. FPGAs (Gandhare and Karthikeyan, 2019) are a group of reconfigurable devices allowing for implementing a variety of digital circuits without the need for fabricating a bespoke ASIC. Initially designed to prototype ASICs prior to tape-out, nowadays the FPGAs are used in a plethora of different fields from low-volume electronics (Juan Jose Rodriguez Andina et al., 2017), through space exploration (Jacobs et al., 2012), to large data centers (Weerasinghe et al., 2015). A hardware designer can program an FPGA to act as a central processing unit (CPU) (Calderón et al., 2005), a Graphics Processing Unit (GPU) (Andryc et al., 2013), or a neuromorphic device. FPGAs have shown a lot of promise in accelerating neuromorphic workloads, resulting in significant performance gains over CPU- and GPU-based solutions (Podobas and Matsuoka, 2017; Ju et al., 2020), with examples of even outperforming ASIC-based neuromorphic systems (e.g., for brain simulation (Lindqvist and Podobas, 2024)).
@@
1.1. Motivation
There are many excellent surveys on neuromorphic systems and spiking neural networks (SNNs) from different perspectives, tied to hardware (Schuman et al., 2017), specific devices (Davies et al., 2021) (e.g., Intel Loihi), different models (Shrestha et al., 2022), and more. For the prior work surveying FPGAs and SNNs (Maślanka and Gorgoń, 2012; Mehrabi and van Schaik, 2024; Maguire et al., 2007), we found that they are very narrow in scope or outdated and that they do not cover a number of significant areas, including the analysis of trends and projections of where the field would be in the years to come. Hence, this survey is – to the best of our knowledge – the most complete review of neuromorphic hardware architectures on FPGAs and includes novel analysis of trends and projections that are not found in any other body of work on the field.
1.2. Method
@@
    ( TITLE-ABS-KEY ( "spiking model*" OR "spiking neur* model*" OR "spiking network*" OR "Spiking Neur* Network*" OR neuromorphic OR ( ( biological* OR realistic* ) W/3 ( "Neur* Network*" OR "neur* model*" ) ) ) AND TITLE-ABS-KEY ( "coarse grained reconfigurable architecture*" OR "coarse grained reconfigurable array*" OR cgra OR fpga OR "field programmable gate array*" OR "reconfigurable architecture*" ) )
  2. (2)
@@ 2. Background
2.1. Field-Programmable Gate Arrays
@@
Figure 1. Abstract view of an FPGA, showing reconfigurable logic (blue), DSP-blocks (green), BRAM (red), and hardened memory controller and processor blocks. Illustration based on (Kuon et al., 2008).
[…]
Diagrams representing the Taxonomy classes’ main features.
@@
Field-Programmable Gate Arrays (FPGAs)(Kuon et al., 2008) belong to a family of reconfigurable architectures, whose underlying silicon can be reprogrammed to yield different functionality. Historically used to prototype ASICs before tape-out, nowadays the FPGAs are used in various different fields from low-volume electronics, through embedded systems and telecommunications to high-performance computing (HPC)(Podobas et al., 2017; Zohouri et al., 2016). Performance of the FPGA-based systems can generally be better than that of the systems based on CPUs or GPUs but worse than that of custom ASICs. Another strength of using FPGAs is that they can be iteratively tailored to the needs of the application without the cost of developing a new ASIC every time. To put this matter into perspective, an FPGA capable of executing tens of billions of floating-point operations per second (tens of TFLOPs) can reach the price of thousands of dollars, while creating a similarly capable ASIC is between two and three orders of magnitude more expensive. Fundamentally, every FPGA consists of a number of logic blocks (also called configurable logic blocks or logic elements), a system of programmable interconnects (called fabric) that routes signals between the logic blocks and input/output (I/O) blocks for communication with external devices. Logic blocks are usually made up of a look-up table (LUT) that stores predefined list of logic outputs for any combination of logic input vectors and standard logic circuitry, such as multiplexers, full adders and flip-flops. The content of these LUTs can be programmed, and by connecting several LUTs together, we can achieve any digital functionality. Originally, FPGAs comprised only the aforementioned logic blocks and fabric, but it was soon realized that certain digital circuits (e.g., multipliers) consumed much of the FPGA’s logic resources. Thus, dedicated (called hardened) functionality was added to the subsequent iterations of the FPGA chips, e.g., bespoke on-chip memory called block random access memory (BRAM), digital signal processing (DSP) blocks (which, in some FPGAs, include floating-point units) and many others (phased locked loops, PCIe controllers, etc.). Figure 1 shows a view of an FPGA.
@@
There are multiple ways to program FPGAs. The most well-known way that also gives the best results – but also the one with the steepest learning curve – is to design logic at the register transfer level (RTL) using hardware description languages (HDLs) such as VHDL or Verilog. Another approach, which has become increasingly popular lately, is to write the intended functionality using programming languages such as C or C++ and let the compilers automatically transform the code into hardware. This is called High-Level Synthesis (HLS)(Nane et al., 2015) and allows engineers and researchers without background in electrical engineering or computer science to work with FPGAs.
2.2. Spiking Neural Networks
@@
Biological neurons communicate via action potentials - voltage pulses traveling from one neuron to another, caused by rapid changes in neurons’ membrane potentials - commonly referred to as spikes or spike events. Networks built using neurons that mimic this behavior are called Spiking Neural Networks (SNNs). SNNs are often called the third generation networks(Maass, 1997) that add a notion of time to the artificial neural networks (ANNs). SNNs have been used as tools to understand the biological brain in the neuroscience community (e.g., to study cortical columns(Potjans and Diesmann, 2014)) through simulators such as Neuron(Hines and Carnevale, 1997), NEST(Gewaltig and Diesmann, 2007), or Brian(Stimberg et al., 2019). Lately, these models have also been considered for solving machine learning (ML) problems. Moreover, SNNs are believed to reduce the energy consumption needed to solve those problems and/or solve them faster, primarily by leveraging the sparseness of spike communication. Spiking neurons can be arranged in various topologies that fit different scenarios, from classification tasks to neuroscientific experiments. Those include, but are not limited to: Feed-Forward (FF) - especially the fully-connected subtype (FF-FC) which is the spiking equivalent of the multilevel perceptron (MLP), Spiking Convolutional Neural Networks (SCNNs) - which despite the feed-forward operation deserve a separate category, because of its popularity and utilization of convolutional kernels and not direct connections between neurons, Liquid State Machines (LSM)(Maass, 2011) - including a randomly-connected reservoir of neurons and a fully connected output layer, randomly-connected (RAND) - being a reservoir of recurrently- and randomly-connected neurons, all-to-all-connected (A2A) and biological structures represented as SNNs [^1] Presented abbreviations are not necessarily used unanimously. Still, they commonly appear in the literature. - e.g., Central Pattern Generators (CPGs)(Wilson, 1961; Jankowska and Roberts, 1972), which are neuronal circuits that when activated can produce rhythmic motor patterns such as walking or breathing. As for the communication, the most popular way of exchanging spike events is the Address Event Representation (AER)(Mahowald, 1994) protocol, representing the event as the destination address of a neuron or a processing element (PE) responsible for computing its dynamics, and the timestep when it occurred.
2.2.1. Neuron models
@@
Table 1. The most commonly used spiking neuron models in our survey (see Section 4 for analysis). The biological plausibility and complexity of the model metrics are based on (Izhikevich, 2004) and our own analysis. The complexity metric relates to the estimated number of floating-point operations (addition, multiplication, etc.) necessary to advance to the next timestep (1 ms) with a selected neuron model ($v$ membrane potential, $I$ synaptic current, $R$ leakage resistance, $g_{L}$ leakage conductance, $C$ membrane capacitance, $g_{K}$ conductance of the potassium ion channel, $g_{NA}$ conductance of the sodium ion channel; other parameters are model-specific, and interested readers are referred to the related literature for detailed explanations).
Neuron Models
Name Formula Bio. plausibility Complexity
Integrate-and-Fire (IF) $\begin{array}[]{ll}\frac{dv}{dt}=RI\end{array}$ V. Low  5 FLOPs
Leaky IF (LIF) $\begin{array}[]{ll}\frac{dv}{dt}=-(v-V_{reset})+RI\end{array}$ Low  10 FLOPs
Izhikevich (IZH) $\begin{cases}\frac{dv}{dt}=0.04v^{2}+5v+140-u+I\\ \frac{du}{dt}=a(bv-u)\end{cases}$ Low  60 FLOPs
FitzHugh-Nagumo (FHN) $\begin{cases}\frac{dv}{dt}=v-\frac{v^{3}}{3}-w+I\\ \frac{dw}{dt}=0.08(v+0.7-0.8w)\end{cases}$ Medium  100 FLOPs
Adaptive IF (AdEX) $\begin{cases}\frac{dv}{dt}=\frac{-g_{L}(v-E_{L})+g_{L}\Delta_{T}exp(\frac{v-v_{T}}{\Delta_{T}})-w+I}{C}\\ \frac{dw}{dt}=a(v-E_{L})-w\end{cases}$ Medium  200 FLOPs
Hodgkin-Huxley (HH) $\begin{cases}\frac{du}{dt}=\\ \frac{-g_{K}n^{4}(U-V_{K})-g_{Na}m^{3}h(U-V_{Na})-g_{leak}(U-u_{leak})+I}{C_{m}}\\ \frac{dn}{dt}=\alpha_{n}(U)(1-n)-\beta_{n}(U)n\\ \frac{dm}{dt}=\alpha_{m}(U)(1-m)-\beta_{m}(U)m\\ \frac{dh}{dt}=\alpha_{h}(U)(1-h)-\beta_{h}(U)h\end{cases}$ High  1200 FLOPs
@@
An SNN uses a spiking neuron model to describe the functionality of generating outgoing spikes. Those models are formulated using a set of ordinary differential equations (ODEs) that describe how a neuron cell membrane potential interacts with an injected current or modified conductance over time. In addition to the ODEs, models have a condition for emitting a spike, typically when the membrane voltage potential reaches a certain value, after which the neuron enters a state of inactivity (called refractory period). Neuron models vary in biological plausibility and complexity, with different models suitable for different use cases. Table 1 outlines six important neuron models, their mathematical description, biological plausibility, and an estimation of how many floating-point operations (FLOPs) are needed to advance one millisecond of simulation time using them. The Integrate-and-Fire (IF) neuron model, which is simple and biologically implausible, is also the least expensive to compute and requires only a handful of operations. On the other hand, the Nobel prize-winning Hodgkin-Huxley (HH) model is complex, describes in high detail the behavior of biological neurons, but it requires thousands of operations to solve the ODEs in a numerically stable manner. Hence, the choice of neuron model comes primarily from the use case: an SNN used for detailed simulation of the biological system might need HH-like neurons (e.g., the OpenWorm(Sarma et al., 2018) project is based on these), whereas a machine learning system performing image inference might not need that level of realism, and could use a less expensive model (e.g., the Leaky Integrate-and-Fire (LIF) model, which is a popular choice for image inference(Diehl and Cook, 2015)). While the neurons in Table 1 were selected based on them being the most popular in neuromorphic hardware (as we will see in Sections 3 and 4), there are multiple other models worth mentioning, each positioned somewhere in the span of complexity between LIF and HH(Hodgkin and Huxley, 1952). These include the Spike Response Model (SRM)(Gerstner et al., 2014), pulsed neurons(Waldemark et al., 1998), Message-Based Event-Driven (MBED) neurons(Claverol, 2000) and Traub’s model(Traub et al., 1991), to name a few.
2.2.2. Synapse Models

### v3

@@ Abstract
1 Introduction
@@
Computers and their performance improvement have historically relied on two pillars: Moore’s law [1] and Dennard’s scaling [2]. The former was formulated by Gordon Moore in the 1960s and states that the number of transistors per unit area double every 24 months, primarily due to improvements in lithography techniques. The latter was introduced by Richard Dennard, who found that as transistors grow smaller, the energy consumption reduces, resulting in constant power density. These two observations worked in tandem, allowing us to build more complex high-performing computers (Moore’s law) without them consuming needlessly large amounts of energy (Dennard’s scaling). Unfortunately, Dennard’s scaling ended in 2004-2005 [3], and we are currently witnessing the slow-down of Moore’s law [1] as well as the emergence of Dark Silicon [4] without any in-place replacement technology at hand [5]. Instead, researchers and industry are looking for alternative ways of computing in hope of continuing the trend in computer performance we have seen so far. These technologies are collectively called post-Moore technologies [6, 5] and include quantum computers [7], adiabatic reversible logic [8], and analog computing [9]. Two post-Moore technologies, however, are perhaps the most salient alternatives: Neuromorphic– [10] and Reconfigurable–computing [11, 12].
@@
Introduced in the 1980s by Carver Mead, neuromorphic systems [13] are computers built to replicate – to various extent – the impressive computing capabilities of the biological brain. Initially designed to replicate brain circuitry using analog electronic components, neuromorphic systems today are implemented in different technologies, ranging from digitally synchronous (e.g., Intel Loihi [14] or IBM TrueNorth [15]), mixed analog/digital (e.g., BrainDrop [16]), fully analog (e.g. [17]), to even memristor- or photonics-based systems [18, 19]. Regardless of how they are implemented, they share two fundamental traits: (i) they communicate through discrete events called spikes and (ii) they are programmed in a non-imperative way, generally by encoding the problem to be solved through a spiking neural network (SNN) [20]. Solving a problem using a neuromorphic system, often results in a significantly more energy-efficient solution in several application domains [21] as opposed to a traditional von Neumann system [22]. A large number of functionalities of neuromorphic systems are driven by discoveries and innovation in neuroscience (e.g., Spike-Timing Dependent Plasticity (STDP) [23], Bayesian Confidence Propagation Neural Network (BCPNN) [24], or Eligibility Propagation (E-prop) [25]), which has motivated many researchers to abandon non-reconfigurable and expensive application-specific integrated circuits (ASICs) and leverage the reconfigurable Field-Programmable Gate Arrays (FPGAs) to keep up with the fast pace of neuroscientific research. FPGAs [26] are a group of reconfigurable devices allowing for implementing a variety of digital circuits without the need for fabricating a bespoke ASIC. Initially designed to prototype ASICs prior to tape-out, nowadays the FPGAs are used in a plethora of different fields from low-volume electronics [27], through space exploration [28], to large data centers [29]. A hardware designer can program an FPGA to act as a central processing unit (CPU) [30], a Graphics Processing Unit (GPU) [31], or a neuromorphic device. FPGAs have shown a lot of promise in accelerating neuromorphic workloads, resulting in significant performance gains over CPU- and GPU-based solutions [32, 33], with examples of even outperforming ASIC-based neuromorphic systems (e.g., for brain simulation [34]).
@@
1.1 Motivation
There are many excellent surveys on neuromorphic systems and spiking neural networks (SNNs) from different perspectives, tied to hardware [10], specific devices [21] (e.g., Intel Loihi), different models [35], and more. For the prior work surveying FPGAs and SNNs [36, 37, 38], we found that they are very narrow in scope or outdated and that they do not cover a number of significant areas, including the analysis of trends and projections of where the field would be in the years to come. Hence, this survey is – to the best of our knowledge – the most complete review of neuromorphic hardware architectures on FPGAs and includes novel analysis of trends and projections that are not found in any other body of work on the field.
1.2 Method
@@
    ( TITLE-ABS-KEY ( "spiking model*" OR "spiking neur* model*" OR "spiking network*" OR "Spiking Neur* Network*" OR
    neuromorphic OR ( ( biological* OR realistic* ) W/3 ( "Neur* Network*" OR "neur* model*" ) ) )
    AND TITLE-ABS-KEY ( "coarse grained reconfigurable
    architecture*" OR "coarse grained reconfigurable array*" OR cgra OR fpga OR "field programmable gate array*" OR
    "reconfigurable architecture*" ) )
  2. 2.
@@ 2 Background
2.1 Field-Programmable Gate Arrays
@@
Figure 1: Abstract view of an FPGA, showing reconfigurable logic (blue), DSP-blocks (green), BRAM (red), and hardened memory controller and processor blocks. Illustration based on [12].
@@
Field-Programmable Gate Arrays (FPGAs)[12] belong to a family of reconfigurable architectures, whose underlying silicon can be reprogrammed to yield different functionality. Historically used to prototype ASICs before tape-out, nowadays the FPGAs are used in various different fields from low-volume electronics, through embedded systems and telecommunications to high-performance computing (HPC)[39, 40]. Performance of the FPGA-based systems can generally be better than that of the systems based on CPUs or GPUs but worse than that of custom ASICs. Another strength of using FPGAs is that they can be iteratively tailored to the needs of the application without the cost of developing a new ASIC every time. To put this matter into perspective, an FPGA capable of executing tens of billions of floating-point operations per second (tens of TFLOPs) can reach the price of thousands of dollars, while creating a similarly capable ASIC is between two and three orders of magnitude more expensive. Fundamentally, every FPGA consists of a number of logic blocks (also called configurable logic blocks or logic elements), a system of programmable interconnects (called fabric) that routes signals between the logic blocks and input/output (I/O) blocks for communication with external devices. Logic blocks are usually made up of a look-up table (LUT) that stores predefined list of logic outputs for any combination of logic input vectors and standard logic circuitry, such as multiplexers, full adders and flip-flops. The content of these LUTs can be programmed, and by connecting several LUTs together, we can achieve any digital functionality. Originally, FPGAs comprised only the aforementioned logic blocks and fabric, but it was soon realized that certain digital circuits (e.g., multipliers) consumed much of the FPGA’s logic resources. Thus, dedicated (called hardened) functionality was added to the subsequent iterations of the FPGA chips, e.g., bespoke on-chip memory called block random access memory (BRAM), digital signal processing (DSP) blocks (which, in some FPGAs, include floating-point units) and many others (phased locked loops, PCIe controllers, etc.). Figure 1 shows a view of an FPGA.
@@
There are multiple ways to program FPGAs. The most well-known way that also gives the best results – but also the one with the steepest learning curve – is to design logic at the register transfer level (RTL) using hardware description languages (HDLs) such as VHDL or Verilog. Another approach, which has become increasingly popular lately, is to write the intended functionality using programming languages such as C or C++ and let the compilers automatically transform the code into hardware. This is called High-Level Synthesis (HLS)[41] and allows engineers and researchers without background in electrical engineering or computer science to work with FPGAs.
2.2 Spiking Neural Networks
@@
Biological neurons communicate via action potentials - voltage pulses traveling from one neuron to another, caused by rapid changes in neurons’ membrane potentials - commonly referred to as spikes or spike events. Networks built using neurons that mimic this behavior are called Spiking Neural Networks (SNNs). SNNs are often called the third generation networks[20] that add a notion of time to the artificial neural networks (ANNs). SNNs have been used as tools to understand the biological brain in the neuroscience community (e.g., to study cortical columns[42]) through simulators such as Neuron[43], NEST[44], or Brian[45]. Lately, these models have also been considered for solving machine learning (ML) problems. Moreover, SNNs are believed to reduce the energy consumption needed to solve those problems and/or solve them faster, primarily by leveraging the sparseness of spike communication. Spiking neurons can be arranged in various topologies that fit different scenarios, from classification tasks to neuroscientific experiments. Those include, but are not limited to: Feed-Forward (FF) - especially the fully-connected subtype (FF-FC) which is the spiking equivalent of the multilevel perceptron (MLP), Spiking Convolutional Neural Networks (SCNNs) - which despite the feed-forward operation deserve a separate category, because of its popularity and utilization of convolutional kernels and not direct connections between neurons, Liquid State Machines (LSM)[46] - including a randomly-connected reservoir of neurons and a fully connected output layer, randomly-connected (RAND) - being a reservoir of recurrently- and randomly-connected neurons, all-to-all-connected (A2A) and biological structures represented as SNNs [^1] Presented abbreviations are not necessarily used unanimously. Still, they commonly appear in the literature. - e.g., Central Pattern Generators (CPGs)[47, 48], which are neuronal circuits that when activated can produce rhythmic motor patterns such as walking or breathing. As for the communication, the most popular way of exchanging spike events is the Address Event Representation (AER)[49] protocol, representing the event as the destination address of a neuron or a processing element (PE) responsible for computing its dynamics, and the timestep when it occurred.
2.2.1 Neuron models
{adjustwidth}
-1.6cm Neuron Models Name Formula Bio. plausibility Complexity Integrate-and-Fire (IF) $\begin{array}[]{ll}\frac{dv}{dt}=RI\end{array}$ V. Low  5 FLOPs Leaky IF (LIF) $\begin{array}[]{ll}\frac{dv}{dt}=-(v-V_{reset})+RI\end{array}$ Low  10 FLOPs Izhikevich (IZH) $\begin{cases}\frac{dv}{dt}=0.04v^{2}+5v+140-u+I\\ \frac{du}{dt}=a(bv-u)\end{cases}$ Low  60 FLOPs FitzHugh-Nagumo (FHN) $\begin{cases}\frac{dv}{dt}=v-\frac{v^{3}}{3}-w+I\\ \frac{dw}{dt}=0.08(v+0.7-0.8w)\end{cases}$ Medium  100 FLOPs Adaptive IF (AdEX) $\begin{cases}\frac{dv}{dt}=\frac{-g_{L}(v-E_{L})+g_{L}\Delta_{T}exp(\frac{v-v_{T}}{\Delta_{T}})-w+I}{C}\\ \frac{dw}{dt}=a(v-E_{L})-w\end{cases}$ Medium  200 FLOPs Hodgkin-Huxley (HH) $\begin{cases}\frac{du}{dt}=\\ \frac{-g_{K}n^{4}(U-V_{K})-g_{Na}m^{3}h(U-V_{Na})-g_{leak}(U-u_{leak})+I}{C_{m}}\\ \frac{dn}{dt}=\alpha_{n}(U)(1-n)-\beta_{n}(U)n\\ \frac{dm}{dt}=\alpha_{m}(U)(1-m)-\beta_{m}(U)m\\ \frac{dh}{dt}=\alpha_{h}(U)(1-h)-\beta_{h}(U)h\end{cases}$ High  1200 FLOPs
@@ 2.2.1 Neuron models
Table 1: The most commonly used spiking neuron models in our survey (see Section 4 for analysis). The biological plausibility and complexity of the model metrics are based on [50] and our own analysis. The complexity metric relates to the estimated number of floating-point operations (addition, multiplication, etc.) necessary to advance to the next timestep (1 ms) with a selected neuron model ($v$ membrane potential, $I$ synaptic current, $R$ leakage resistance, $g_{L}$ leakage conductance, $C$ membrane capacitance, $g_{K}$ conductance of the potassium ion channel, $g_{NA}$ conductance of the sodium ion channel; other parameters are model-specific, and interested readers are referred to the related literature for detailed explanations).
@@
An SNN uses a spiking neuron model to describe the functionality of generating outgoing spikes. Those models are formulated using a set of ordinary differential equations (ODEs) that describe how a neuron cell membrane potential interacts with an injected current or modified conductance over time. In addition to the ODEs, models have a condition for emitting a spike, typically when the membrane voltage potential reaches a certain value, after which the neuron enters a state of inactivity (called refractory period). Neuron models vary in biological plausibility and complexity, with different models suitable for different use cases. Table 1 outlines six important neuron models, their mathematical description, biological plausibility, and an estimation of how many floating-point operations (FLOPs) are needed to advance one millisecond of simulation time using them. The Integrate-and-Fire (IF) neuron model, which is simple and biologically implausible, is also the least expensive to compute and requires only a handful of operations. On the other hand, the Nobel prize-winning Hodgkin-Huxley (HH) model is complex, describes in high detail the behavior of biological neurons, but it requires thousands of operations to solve the ODEs in a numerically stable manner. Hence, the choice of neuron model comes primarily from the use case: an SNN used for detailed simulation of the biological system might need HH-like neurons (e.g., the OpenWorm[51] project is based on these), whereas a machine learning system performing image inference might not need that level of realism, and could use a less expensive model (e.g., the Leaky Integrate-and-Fire (LIF) model, which is a popular choice for image inference[52]). While the neurons in Table 1 were selected based on them being the most popular in neuromorphic hardware (as we will see in Sections 3 and 4), there are multiple other models worth mentioning, each positioned somewhere in the span of complexity between LIF and HH[53]. These include the Spike Response Model (SRM)[54], pulsed neurons[55], Message-Based Event-Driven (MBED) neurons[56] and Traub’s model[57], to name a few.
2.2.2 Synapse Models